Please help me solve this problem.


A fair die is rolled. Find the probability of obtaining a number greater than 3.

Found 2 solutions by stanbon, subudear:
Answer by stanbon(75887)

Please help me solve this problem.

A fair die is rolled. Find the probability of obtaining a number greater than 3.
-----
P(4 or 5 or 6) = 3/6 = 1/2
================================
Cheers,
Stan H.

Answer by subudear(62)

numbers greater then 3 on a die are 4, 5 ,6
so probability of obtaining a number greater than 3 is 3/6 = 1/2